OpenStudy (rissyroo118):

which equation can be used to find the two numbers whose ratio is 3 to 2 and that have a sum of 35? help?? 3\2x=35 2\3x=35 3x – 2x = 35 3x + 2x = 35

OpenStudy (anonymous):

You have to use x/y = 3/2 ---- (1) and x + y = 35 ----(2) and then solve the two equations.